OpenStudy (anonymous):

What is the sum of the square root of -2 and the square root of -18?

OpenStudy (anonymous):

square root of 2 and -2 are same so value is 1.41

OpenStudy (anonymous):

value of square root of -8 = 4.24 which is the value of square root of 8

OpenStudy (anonymous):

so add up the values and tell me the answer

OpenStudy (anonymous):

I got 5.65

OpenStudy (anonymous):

yeahhh

OpenStudy (anonymous):

That's not one of my choices though

OpenStudy (anonymous):

My choices are...

OpenStudy (anonymous):

they are ?

OpenStudy (anonymous):

\[4\sqrt{2}\]

OpenStudy (anonymous):

\[4\sqrt{2i}\]

OpenStudy (anonymous):

\[5\sqrt{2}\]

OpenStudy (anonymous):

and lastly \[5\sqrt{2i}\]

OpenStudy (anonymous):

okay so \[3\sqrt{2} = 18 do u know that\]

OpenStudy (anonymous):

:P lol weir hand writitng hahahhaahha

OpenStudy (anonymous):

okay srs do u know that

OpenStudy (anonymous):

No Idon't understand any of this

OpenStudy (anonymous):

okay so 18 = 2 * 3 * 3 u take 3 out of the root as there are 2 3s remember in square root u need to take away a number that comes twice as 1 like if there is 2 and 2 then take it as single 2